### v1

# Calibration of the strain amplitude recorded with DAS using a strainmeter array

Abstract
The power of distributed acoustic sensing (DAS) lies in its ability to sample deformation signals along an optical fiber at hundreds of locations with only one interrogation unit (IU). While the IU is calibrated to record ‘fiber strain’, the properties of the cable and its coupling to the rock control the ‘strain transfer rate’ and hence how much of ‘rock strain’ is represented in the recorded signal. We use DAS recordings in an underground installation colocated with an array of strainmeters in order to calibrate the ‘strain transfer rate’ in situ, using earthquake signals between 0.05 Hz and 0.1 Hz. A tight-buffered cable and a standard loose-tube telecommunication cable (running in parallel) are used, where a section of both cables loaded down by loose sand and sand bags is compared to a section, where cables are just unreeled on the floor. The ‘strain transfer rate’ varies between 0.13 and 0.53 depending on cable and installation type. The sandbags show no obvious effect and the tight-buffered cable generally provides a larger ‘strain transfer rate’. Calibration of the ‘strain transfer rate’ with respect to the strainmeter does not depend on wave propagation parameters. Hence it is applicable to the large amplitude surface wave signal in a strain component almost perpendicular to the great-circle direction for which a waveform comparison with seismometer data does not work. The noise background for ‘rock strain’ in the investigated band is found at about an rms-amplitude of 0.1 nstrain in 1/6 decade for the tight-buffered cable. This allows a detection of marine microseisms at times of high microseism amplitude.
1 Introduction
Distributed Acoustic Sensing (DAS) measures dynamic strain in an optical fiber. The DAS interrogation unit (IU) sends laser-generated coherent light into an optical fiber possibly extending for tens of kilometers. The light gets partly back-scattered by Rayleigh scattering due to manufacturing imperfections along the optical fiber and is sensitive to rapid variations in strain, commonly referred to as dynamic strain. Pioneering work by Dakin (1990), followed by Taylor and Lee (1993), recognized the potential of coherent Rayleigh back-scatter in assessing spatial disturbances in optical fibers. Following this work a range of techniques has been developed and implemented in IUs to measure the phase of the back-scattered signal. Hartog (2017) summarizes these progressions and outlines applications of distributed optical fiber sensors across various domains.
In seismology, the power of DAS lies in its ability to sample deformation signals along an optical fiber at hundreds of locations over distances of many kilometers with a single IU. In some applications, unused telecommunication infrastructures (so-called dark-fibers) can be leveraged, which significantly reduces the necessary effort for field deployment. Lindsey and Martin (2021) and Li et al. (2021) present an overview of fields of application in geosciences. Most of these applications rely only on the phase information in the recorded data.
Use cases, which rely on amplitude are less frequent. Such a use case would be the measurement of volume strain for the reduction of Newtonian Noise (Harms, 2015; Harms et al., 2022), as being planned for the Einstein Telescope (ET Steering Committee, 2020).
With well calibrated DAS strain recordings on cables running in six different directions from a single location, the full strain tensor could be composed. This requires a proper calibration of the ‘strain transfer rate’ (the fraction of ‘rock strain’ picked up as ‘fiber strain’) and a sufficiently low detection level, such that the strain background signal is resolved. Rademacher (2024) recently pointed at the limitations which still exist for DAS in both respects. We investigate these in the current study.
Signals recorded by the DAS IU represent the deformation of the fiber (denoted here as ‘fiber strain’). However, observations showing that the signal from two colocated cables might differ in amplitude (Azzola et al., 2022, and the current study) suggest that ‘fiber strain’ does not always represent the actual ‘rock strain’. This discrepancy can arise from the coupling of the fiber to the ground through the various layers of the cable.
Reinsch et al. (2017), for example, focus on the internal structure of an optic fiber cable and discuss this discrepancy based on a physical model (Li et al., 2006) of a DAS cable, based on actual material properties. They derive the ‘strain transfer rate’ as the ratio between ‘fiber strain’ and ‘rock strain’. In particular gel layers in loose-tube cables let the ‘strain transfer rate’ be less than 1 due to their small value of Young’s modulus. These thixotropic fluids are used to protect the fibers from damage. Reinsch et al. (2017) estimate that during measurement their yield point is not exceeded, such that they behave in a linear elastic way. In practice the parameters of the different layers in the cable are not available from manufacturers and the coupling to the rock is controlled by the actual installation conditions. For this reason the ‘strain transfer rate’ can be determined empirically only.
Comparisons of DAS recordings with other types of observations are often carried out with respect to colocated seismometers. This approach involves translating strain or strain rate into displacement or velocity (or vice versa). This is only possible for non-dispersive plane waves with known ray-parameter. In this study, we directly compare the recordings of a DAS IU with those from the long-running array of Invar-wire strainmeters at the Black Forest Observatory (BFO). Both, the DAS recorded strain and the signal from the strainmeter can be expected to represent the same rock deformation, independent of its cause and independent of wave parameters in particular. Hence, this comparison allows for direct calibration of the DAS system’s amplitude response. Based on this calibration we estimate the detection threshold for strain recordings. The signal-to-noise ratio of both systems is appropriate for this purpose in the frequency band between 0.05 Hz and 0.1 Hz.
2 DAS installation
and instruments
Interrogation unit (IU) Febus A1-R | |
Software: | version 2.2.2 |
optical wavelength: | 1550.12 nm |
Fiber length: | 2200 m |
Pulse width: | 5 m |
Block rate: | 1 Hz |
Pulse rate frequency: | 5 kHz |
Ampli power: | 28 dBm |
Sampling resolution: | 80 cm |
Gauge length (GL): | 50 m |
Derivation time (DT): | 20 ms |
Channel spacing: | 24.8 m |
Black Forest Observatory (BFO) is situated in a former silver mine in the central Black Forest, Germany (Emter et al., 1994). The gallery that hosts the instruments (Fig. 1) is mostly horizontal and is excavated in granite. The granite is covered with Triassic sedimentary rocks (Emter et al., 1994). The study focuses on two straight sections of the tunnel, ‘Anton Gang’ and ‘Vorstollen’. Overburden increases with distance to the tunnel entrance, is about 100 m at the ‘Anton Gang’ and reaches 170 m at the ‘Strainmeter array’.
2.1 Distributed acoustic sensing (DAS)
Two fiber optic cables are installed in the tunnel. They are unreeled on the floor of the gallery and are in direct contact with the formation. In the section called ‘Anton Gang’ the cables are loaded down by sand and tightly spaced sandbags in order to improve the mechanical coupling to the rock, provide thermal shielding and cover the cable against water dripping from the gallery ceiling. The blue cable in Fig. 1 is a standard flexible telecommunication cable with loose-tube fibers embedded in gel. The green cable in Fig. 1 has a stiff jacket containing tight-buffered fibers (see section S1.2 in the supplemental material for additional details). We use two single mode fibers in each of the cables and splice them in series, such that the ‘Anton Gang’ and the ‘Vorstollen’ are both sampled four times. The signals from the fibers are sampled by a Febus A1-R interrogation unit (IU) installed in the laboratory (Fig. 1). The DAS recording parameters are listed in Table 1.
The Febus A1-R IU operates on the principle of differential phase-measuring distributed acoustic sensing. The light that is back-scattered in the optical fiber is mixed with a reference signal to measure the differential phase over a gauge length, subsequently leading to distributed strain-rate measurements, which is converted to strain by integration over time. The A1-R falls within the category of heterodyne Distributed Vibration Sensing (hDVS) systems, using a single-pulse heterodyne approach for phase detection, as described by Pan et al. (2011).
2.2 Reference instruments
BFO operates an array of three well calibrated, 10 m long, horizontal, Invar-wire strainmeters (Fig. 1). The SEED codes for the instruments are II.BFO.00.BSA, II.BFO.00.BSB, and II.BFO.00.BSC. Their design is based on the instruments by King and Bilham (1976) and are discussed in more detail by Zürn et al. (2015). Agnew (1986) and Zürn (2012) discuss instruments of this type and their properties. These instruments are primarily designed to record very-long period signals, such as tidal strain or Earth’s free oscillations. At short signal period they show a linear parasitic sensitivity to vertical ground acceleration because of the inertia of the pick-up system. They provide useful strain signals at frequencies below 1 Hz. We use the instruments at frequencies below 0.1 Hz (which provides a safety-margin to the parasitic response) as a reference to present ‘rock strain’.
We use signals recorded by the STS-2 broad-band seismometer (Fig. 1) to demonstrate the limitations of waveform comparisons with particle velocity derived strain. The SEED codes for its three components are GR.BFO..BHZ, GR.BFO..BHN, and GR.BFO..BHE.
3 Data and data processing
3.1 Available data
The installation of DAS cables as shown in Fig. 1 was used for recording from May 22, 2022 until March 13, 2023. The IU was moved on February 7, 2023 from the laboratory to the electronics vault in the mine (see Fig. 1 and Figs. S1 and S2 in the supplemental material for additional details). In both, ‘Anton Gang’ and ‘Vorstollen’, we focus on four almost colocated read-out locations that are separated by 5 m to 15 m. In all cases, the section of gauge length completely falls into the ‘Anton Gang’ or ‘Vorstollen’, such that the strain recordings are representative of the azimuth of these gallery sections.
From the GEOFON catalog (Quinteros et al., 2021) we find 84 earthquakes with moment magnitude larger than 6 in the recording time period. 21 of them show a maximum strain amplitude of larger than 1 nstrain in a visual inspection, which is considered large enough to provide a sufficient signal-to-noise ratio in the DAS data. For 19 of the events DAS data is available and allows an analysis. Event characteristics are detailed in the supplementary material (see Table S2). The largest amplitudes are found in the surface wave train. For body-waves, the horizontal strain amplitude is the smaller the steeper the ray incidence.
The strongest signals are recorded from the main shocks (Mw 7.7 and Mw 7.6) of the Kahramanmaraş earthquake sequence (Melgar et al., 2023) on February 6th 2023. Their body-wave signals have large enough amplitude of a few nstrain to be included in the analysis. This is due to the larger earthquake magnitude and the smaller epicentral distance (), compared to the other analyzed events in this study, which both result in rather large amplitude in general. In the set of analyzed events their surface waves provide by far the largest strain amplitudes of about 130 nstrain and about 250 nstrain, respectively, as measured by the Invar-wire strainmeters in the investigated frequency band of 0.05 Hz to 0.1 Hz.
For the comparison of strain waveforms, we can use 46 recordings for each combination of cable (tight-buffered and loose-tube) and location (‘Anton Gang’ and ‘Vorstollen’). For the loose-tube cable in ‘Anton Gang’ we use only 44 recordings, as the configuration was modified on 2023, February 7th.
3.2 Pre-processing of DAS data
Signals sampled simultaneously on a section of the tight-buffered cable rolled up on a ‘Drum’ (see Fig. 1 and Table S1 in the supplemental material) get stacked and subtracted from the recordings in order to reduce common mode laser noise. For a gauge-length of 50 m this noise component clearly dominates the background level of the DAS signal at frequencies below 0.5 Hz and peaks at about 0.07 Hz with ‘fiber strain’-amplitudes of a few nstrain. By subtracting this coherent signal component, the background level is lowered by up to 20 dB near 0.1 Hz. Data is converted from strain rate to strain for which the noise floor rapidly increases at low frequencies, such that we focus our analysis on frequencies higher than 0.05 Hz.
3.3 Linear strain from the
strainmeter array
In order to compare the DAS recordings with recordings of the strainmeters, we derive linear strain in either the direction ‘Anton Gang’ or ‘Vorstollen’ by a linear combination of the signals recorded by the three strainmeters. Zürn et al. (2015, their eq. 2) specify linear strain
in azimuth as a function of the components , , and of the 2D strain tensor. Here, ‘Anton Gang’ and ‘Vorstollen’ are in azimuth NE and NE, respectively. Based on this, we compute linear strain in azimuth
from the strain recorded by the three array-instruments, where , , and are in azimuth NE, NE, and NE, respectively. They correspond to the SEED channel names BSA, BSB, and BSC. The rotation matrix
is the inverse of
4 Comparison of
strain measurements
4.1 Strain transfer rate
By fitting the strainmeter data to the DAS data with a linear regression, we compute the ‘strain transfer rate’
where is the DAS time series of ‘fiber strain’ and is the strainmeter time series of ‘rock strain’. Data are filtered by a Butterworth high-pass (0.05 Hz, 4th order) and low-pass (0.1 Hz, 4th order). The average is thus removed from the signals prior to the computation. Fig. 2 (left) shows the ‘strain transfer rate’ and Table 2 summarizes the ranges of values. They are between 0.13 and 0.53 in all cases and primarily depend on location and cable type. The scatter observed in the distribution for one cable at one location is smaller than the differences to the other combinations. The character of scatter of values does not significantly differ between cables or locations. Differences between colocated fibers in the same cable and location are smaller than 0.05 in most cases. We also find no discernible correlation of ‘strain transfer rate’ with signal amplitude, back-azimuth or other earthquake specific parameters (see Figs. S26 and S27 in the supplemental material for example). The largest values of ‘strain transfer rate’ are found for the tight-buffered cable, which was expected. Although the loose-tube cable in the ‘Anton Gang’ is loaded down by sandbags as well, its ‘strain transfer rate’ is the lowest among all fibers. Hence, we do not observe a significant benefit from sand and sandbags to improve overall coupling.
4.2 Waveform similarity
The linear regression used to derive the ‘strain transfer rate’ requires that the ‘fiber strain’ waveform as recorded by DAS is consistent with the ‘rock strain’ waveform obtained from the strainmeters. We use the normalized correlation coefficient
as a measure of waveform similarity. Fig. 3 compares the P-waves radiated by the Mw 7.7 Pazarcık earthquake across all measurement types and illustrates differences in SNR. Fig. 2 (right) shows the distribution of the normalized correlation coefficients obtained for all tested configurations. They primarily depend on signal amplitude and thus on signal-to-noise ratio for the DAS data (see Fig. S31 in the supplemental material). The values are largest and larger than 0.92 for all signals of the Mw 7.7 and Mw 7.6 earthquakes of the Kahramanmaraş earthquake sequence (see Fig. S29 in the supplemental material). For the surface waves recorded by the tight-buffered cable in the ‘Anton Gang’ they are even larger than 0.99. The distribution shown in Fig. 2 (right) indicates a generally better signal quality for data recorded with the tight-buffered cable. The loose-tube cable, if not protected by sand-bags (‘Vorstollen’), appears to be prone to glitches caused by water drops (see Fig. S12 in the supplemental material for example).
The signal-to-noise ratio of the strainmeter generally is better than that of the DAS cables. This can be seen in Fig. 3, where the first P-wave arrival is clearly seen in the seismometer and strainmeter data at 50 s on the time scale. Only after 70 s the signal amplitude is large enough to allow the DAS recording to capture the P-waves. The high signal-to-noise ratio of the ‘rock strain’ allows the strainmeter signal to lock onto the ‘rock strain’-waveform in the DAS data in the linear regression. The high signal-to-noise ratio time series is in the denominator of Eq. (6) and thus there is no correlation between ‘strain transfer rate’ and normalized correlation coefficient found in the results.
4.3 Intercomparison of fibers
at higher frequency
The comparison with strainmeter data is limited to frequencies below 0.1 Hz. Due to their short-period signal energy we can use the body-waves of the Kahramanmaraş earthquakes to extend the study to a larger frequency range by comparing the DAS signals of the main shocks (Mw 7.7 and Mw 7.6) with one another. Hence, the regression is computed with respect to recordings of read-out locations on the tight-buffered fibers. Fig. 4 shows the analysis results for a comparison between different fibers (all with respect to one of the tight-buffered fibers). The normalized correlation coefficients (waveform similarity) generally are higher than 0.970 and reach 1.000 for some of the combinations. At frequencies above 1 Hz the signal-to-noise ratio for S-waves quickly drops because of the decrease of S-wave energy with increasing signal frequency. This results in reduced signal-to-noise ratio and thus waveform similarity and regression coefficients are reduced as well. A frequency dependence of regression factors beyond the scatter is not apparent up to 1 Hz.
In the frequency range up to 1 Hz, colocated fibers in the same cable practically pick up the same strain amplitude. Regression coefficients listed in Table 3 scatter about unity, where the scatter is smaller for the tight-buffered cable (smallest in the ‘Vorstollen’ with a range of 0.96 to 1.02) and largest for the loose-tube cable in the ‘Anton Gang’ (0.80 to 1.23). The ratios of ‘strain transfer rate’ given in Table 2 are consistent with the regression factors for colocated cables in the extended frequency range given in Table 3.
cable | location | minimum | median | maximum |
tight-buffered | ‘Anton Gang’ | 0.39 | 0.46 | 0.53 |
tight-buffered | ‘Vorstollen’ | 0.30 | 0.35 | 0.43 |
loose-tube | ‘Vorstollen’ | 0.21 | 0.27 | 0.36 |
loose-tube | ‘Anton Gang’ | 0.13 | 0.19 | 0.25 |
Values are as presented in Fig. 2 (left). Strainmeter signals are fitted to DAS (distributed acoustic sensing) recorded waveforms by linear regression in the frequency band from 0.05 Hz to 0.1 Hz.
DAS with respect to colocated DAS fibers (in the same cable) | |
tight-buffered ‘Vorstollen’ | 0.96 – 1.02 |
tight-buffered ‘Anton Gang’ | 0.92 – 1.08 |
loose-tube ‘Vorstollen’ | 0.89 – 1.11 |
loose-tube ‘Anton Gang’ | 0.80 – 1.23 |
DAS with respect to colocated DAS fibers (in the other cable) | |
loose-tube vs. tight-buffered ‘Vorstollen’ | 0.72 – 0.83 |
loose-tube vs. tight-buffered ‘Anton Gang’ | 0.32 – 0.47 |
The presented values cover the frequency band from 0.05 Hz to 1.0 Hz.
4.4 Comparison with seismometer data
Strainmeter installations recording at seismic frequencies are rare and available only in a few observatories. In the absence of strainmeters many studies use strain signals estimated from particle velocity recorded by seismometers, which appears possible for plane waves of known incidence and phase-velocity. The scaling relation is given by
where and are horizontal particle velocity and linear strain in azimuth of , respectively, is the azimuth of wave propagation, and is the horizontal component of slowness (see section S4.1 in the supplemental material for additional details). The accuracy of the scaled seismometer signal in representing strain depends on the validity of the plane-wave assumption underlying the conversion. While strainmeters and DAS can be expected to record the same quantity, namely linear strain in a given azimuth, independent of the nature of the signal, this is not the case for the seismometer derived strain signals.
The signal derived from the STS-2 recording in Fig. 3 matches the strain signals quite well near the first P-wave arrival (75 s). In later parts (220 s) the P-wave signal of particle velocity obviously provides only half the amplitude of the signals from the strainmeter and the scaled DAS signals.
The surface waves with good signal-to-noise ratio used in the current study not only propagate at varying slowness (dispersion and superposition of Love- and Rayleigh-waves), they also contain significant non-plane wavefield components due to scattering in particular on continental paths. Wielandt (1993) discusses the consequences of the non-plane nature of waves for the spatial derivatives of displacement. The scaling in Eq. (7) lets vanish for being perpendicular to . In cases where comes close to being perpendicular to the actual strain component is dominated by off-great-circle propagation and non-plane wave contributions, which results in a large scatter of regression coefficients (see Fig. S32 in the supplemental material). In hese cases, seismometers cannot reasonably be used to calibrate DAS signals.
Here, we illustrate the limitations of the comparison with seismometer data using the surface waves recorded on February 23rd 2023 after the Mw 6.8 earthquake in the the Tajikistan-Xinjiang Border Region. The great-circle backazimuth for this earthquake is , which is only by off the direction perpendicular to the ‘Anton Gang’ azimuth of . Fig. 5 (top panel) shows the signal comparison. Love-waves (300 s) and Rayleigh-waves (400 s) are inseparably superimposed. Both are dispersive and arrive at different phase velocity with non-plane wavefronts due to scattering on the continental path. Slowness is used for scaling the seismometer signal, which corresponds to the average phase velocity of about 3.6 km s-1, the value of Rayleigh-wave phase velocity at 0.05 Hz in Southern Germany (Friederich and Huang, 1996, their Table 1). The bottom panel of Fig. 5 shows the waveform residuals with respect to the strainmeter signal. The residual amplitude for the tight-buffered cable is about three to five times the noise level. For the seismometer the largest residual amplitude is five times larger and the waveform mismatch renders the seismometer signal unusable for a calibration. The adjustment of a scaling factor would not solve the problem. The seismometer derived amplitude for the first wave group (300 s – 500 s) is larger than that of strainmeter and DAS, while it is too small after 500 s. Additionally, Fig. 5 shows a significant mismatch throughout the wavetrain, while DAS and strainmeter waveforms are rather consistent. In practically all cases examined the normalized correlation coefficient between seismometer signal and DAS is smaller than between strainmeter signal and DAS (see Fig. S33 in the supplemental material). In most cases it is significantly smaller and in some cases it is even negative (anti-correlation).
5 Detection threshold
for ‘rock strain’
The signal of largest amplitude in the natural seismic background are the marine microseisms. This can be a helpful test signal for the investigation of instrumental properties at frequencies of 0.1 Hz to 1 Hz. In applications for Newtonian Noise mitigation the strain background at even smaller amplitude must be resolved.
We analyze the signals at times of large marine microseism amplitude. The rms-amplitude levels are shown in Fig. 6 (top) together with the pair-wise magnitude-squared coherence (Carter et al., 1973, their eq. 2) of signals (bottom). The signal of the BFO strainmeters shows the signature of the primary microseism peak at about 0.07 Hz as well as the secondary peak at 0.14 Hz. The tight-buffered cable appears to detect the secondary microseism peak, while the scaled noise in the loose-tube fibers is above the background ‘rock strain’ level. The coherence between the BFO strainmeter signal and the tight-buffered fibers (at 620.0 m and 1686.4 m), as well as among the tight-buffered fibers themselves, peaks at 0.8 at the frequency of the secondary microseisms, i.e. 0.14 Hz. The coherence for all other signal combinations reaches a maximum of 0.6 or less. The lower envelope of the coherence curves peaks at 0.3 at 0.14 Hz, which may indicate that all signals contain at least a portion of the marine microseisms. Nevertheless the similarity of the waveforms shown in Fig. 7 is not as good as for the earthquake signals, even though a narrow bandpass is applied. This is consistent with the coherence not exceeding a value of 0.8.
The amplitude levels of the marine microseisms at BFO strongly vary with weather conditions over the North Atlantic and are typically lower during northern hemisphere summer111The seasonal variation of marine microseism energy is seen in the analyses presented by the Waveform Quality Center at Lamont-Doherty Earth Observatory: https://www.ldeo.columbia.edu/~ekstrom/Projects/WQC/MONTHLY_HTML/BFO-II.LHZ-00_time.html.. A similar analysis for signals recorded on 2022-07-10 shows the background level of the BFO strainmeter recordings at below 0.02 nstrain with no indication of a microseism peak. At the same time, the amplitude levels of the DAS recordings range between 0.05 nstrain (0.3 Hz) and 0.4 nstrain (0.05 Hz), with the lowest levels observed for the tight-buffered fibers and rms values decreasing with increasing frequency for all fibers. Coherence between any pair of fibers, or between DAS recording and BFO strainmeter, is below 0.25 at these times. Thus, at times of low marine microseism amplitude, the DAS fibers are not able to pick up the background ‘rock strain’ in the discussed frequency band.
6 Conclusions
The waveform similarity of DAS recorded signals with strainmeter recorded signals as well as in an intercomparison of DAS signals is high for large amplitude signals. For the body- and surface waves recorded after the main shocks (Mw 7.7 and Mw 7.6) of the Kahramanmaraş earthquake sequence on February 6th 2023 the normalized correlation coefficient (NCC) with respect to the strainmeter is larger than 0.92 for all DAS fibers in the frequency band from 0.05 Hz to 0.1 Hz. In an intercomparison of colocated DAS fibers for the body-wave signals up to 1 Hz, the NCC is larger than 0.97 and reaches 1.000 in some cases. For smaller signal amplitude the DAS signal-to-noise ratio worsens and the NCC becomes smaller. The signal quality in this respect is better for the tight-buffered cable and worst for the loose-tube cable in the ‘Vorstollen’, where it is not protected by a sand bed against water dripping from the ceiling.
The DAS recorded ‘fiber strain’ at large amplitude hence well represents the waveform and the signal phase of ‘rock strain’. This is, however, not the case for the signal amplitude. The values of ‘strain transfer rate’ (amplitude ratio of ‘fiber strain’ versus ‘rock strain’), which we calibrate by linear regression, cover a range from 0.13 to 0.53 (Table 2). None of the installed cables provides a ‘strain transfer rate’ equal 1, which would make ‘fiber strain’ equal ‘rock strain’. The largest value is obtained for the tight-buffered cable in the ‘Anton Gang’, where cables are loaded down by loose sand and sand bags. While the tight-buffered cable generally provides a larger ‘strain transfer rate’, the smallest values are obtained for the loose-tube cable in the ‘Anton Gang’. We hence find no consistent effect of the sand bags on coupling the fiber to the rock.
The scatter of ‘strain transfer rate’ values for one cable at one location is smaller than the differences between cables and locations. We find no systematic dependency of ‘strain transfer rate’ on backazimuth of the event, nor do we find any other systematic correlation except with respect to cable type and location. The ‘strain transfer rate’ of the cables cannot be specified independently of local installation conditions.
An amplitude correction factor to be applied to DAS recorded strain not only adjusts the signal amplitude, but also amplifies the instrument generated background noise. Consequently the peak of the secondary marine microseisms is only detected by the installation with largest ‘strain transfer rate’ in Table 3. We estimated the detection threshold for the tight-buffered cable in the ‘Anton Gang’ at 0.1 nstrain rms-amplitude in 1/6 decade near 0.1 Hz.
Even if we attribute the ‘strain transfer rate’ to the cable type and installation here, we cannot rule out the possibility that some other effect reduces the recorded ‘fiber strain’. In a next step, we plan to install a testbed of fibers (just core, cladding, coating, and tight-buffer) directly cemented into the floor of the mine. We expect that this would make the DAS recordings more sensitive and would make ‘fiber strain’ equal ‘rock strain’.
Data and Resources
All data were recorded locally at (Black Forest Observatory [BFO], 1971, https://dx.doi.org/10.5880/BFO). Strainmeter and seismometer data are available through data centers at the (Scripps Institution of Oceanography, 1986, https://dx.doi.org/10.7914/SN/II) and the (Federal Institute for Geosciences and Natural Resources, BGR, 1976, https://dx.doi.org/10.25928/MBX6-HR74), respectively. DAS data can be provided upon request. Event parameters were taken from the GEOFON catalog (Quinteros et al., 2021, https://geofon.gfz-potsdam.de/eqinfo/). Data analysis was carried out with ObsPy (The ObsPy Development Team, 2022, https://dx.doi.org/10.5281/zenodo.6327346). We provide a supplemental text file with figures, tables, and math to further support the findings presented in this paper.
Acknowledgments
We are grateful to Gaetan Calbris and Vincent Lanticq from Febus Optics who provided helpful information on the measurement principle used in the Febus A1-R. Peter Duffner, Felix Bögelspacher, and Leon Merkel supported the installation of the experiment. The cable with the tight-buffered fibers (green cable) has been made available for the experiments through the INSIDE project. INSIDE is supported by the German Federal Ministry for Economic Affairs and Climate Action and the Project Management Jülich (PtJ) under the grant agreement number 03EE4008C.
Appendix S1 Fiber Optic Cable setting
S1.1 Location of sensing points and gauges in the mine
Table S1 details the position of the read-out locations identified near the centers of ‘Anton Gang’ and ‘Vorstollen’. The read-out locations shown in the Figs. S1 and S2 are the ones used in the study.
DAS strain is computed over so-called gauges, defined by the gauge length (GL) parameter, which correspond to the length of optical fiber over which optical dephasing is measured by the IU. The extent of the gauge on the cable plays therefore a significant role in the measured strain. The gauge-sections are displayed in detail in Figs. S1 and S2 for this reason. Fig. S2 applies to the modified cable setup, after February 7, 2023. For each read-out location considered in the analysis on ‘Anton Gang’ and ‘Vorstollen’, the physical location of the sensing point is represented as a symbol with the cable offset indicated in the legend, and the extent of the gauge is represented as a colored line.
The read-out locations were selected because of their position, near the center of ‘Anton Gang’ and ‘Vorstollen’. The figures show the small distance between the read-out locations positioned in each section of interest. It shows also that the associated gauges cover these sections with a consistent azimuth.
S1.2 Specifications of optical fibers
The analysis explores the capacities of two fiber optic cables with different
characteristics.
We use on one hand a standard flexible
telecommunication cable with loose-tube fibers embedded in gel (blue cable on the diagrams).
It is distributed by Prysmian Group.
The print on the cable is:
DRAKA UC FIBRE I/O CT LSHF 3.0kN 4 SM2D
G.652.D/SM7A1 BendBright 60011347 15 20506485 009201015.
On the other hand, we use a second cable that has a stiff jacket containing
tight-buffered fibers (green cable in the diagrams).
It is distributed by Solifos AG and specified as a
‘BRUsens DAS & DTS & Communication Hybrid’ cable.
For distributed acoustic sensing, we use two 2.4 mm simplex elements with
tight-buffered single mode fiber with Aramid, glass strain relief and plastic
outer sheath.
The print on the cable is:
BRUSENS ACOUSTIC TEMPERATURE HYBRID BSAH
2SMF+2MMF+10SMF FIBRE OPTIC CABLE 00241904.
‘Anton Gang’ (azimuth NE) | |||
cables covered with sand and sandbags | |||
UTM32N | |||
location / m | cable | easting / m | northing / m |
396.8 | loose-tube | 450102.6 | 5353139.4 |
620.0 | tight-buffered | 450098.3 | 5353146.5 |
1686.4 | tight-buffered | 450101.6 | 5353141.0 |
1909.6 | loose-tube | 450099.3 | 5353144.9 |
‘Vorstollen’ (azimuth NE) | |||
UTM32N | |||
location / m | cable | easting / m | northing / m |
297.6 | loose-tube | 450028.3 | 5353161.0 |
719.2 | tight-buffered | 450017.0 | 5353161.1 |
1587.2 | tight-buffered | 450023.4 | 5353161.1 |
2008.8 | loose-tube | 450022.0 | 5353161.1 |
‘Anton Gang’ (azimuth NE) after 2023-02-07 | |||
UTM32N | |||
location / m | cable | easting / m | northing / m |
124.0 | tight-buffered | 450101.4 | 5353141.4 |
1190.4 | tight-buffered | 450098.5 | 5353146.2 |
1413.6 | loose-tube | 450102.4 | 5353139.7 |
‘Vorstollen’ (azimuth NE) after 2023-02-07 | |||
UTM32N | |||
location / m | cable | easting / m | northing / m |
223.2 | tight-buffered | 450023.0 | 5353161.1 |
1091.2 | tight-buffered | 450017.4 | 5353161.1 |
1512.8 | loose-tube | 450027.9 | 5353161.0 |
2058.4 | loose-tube | 450019.3 | 5353161.1 |
‘Drum’ | |||
location / m | cable | ||
857.5 | tight-buffered, drum start | ||
1150.0 | tight-buffered, drum center | ||
1442.5 | tight-buffered, drum end | ||
‘Drum’ after 2023-02-07 | |||
location / m | cable | ||
367.5 | tight-buffered, drum start | ||
660.0 | tight-buffered, drum center | ||
952.5 | tight-buffered, drum end |
Appendix S2 Analyzed data
S2.1 Earthquakes used in the analysis
Table S2 details the seismic events used in the analysis with the main source parameters, which are taken from the GEOFON catalog. From the GEOFON catalog (Quinteros et al., 2021) we find 84 earthquakes with moment magnitude larger than 6 in the recording time period. For 21 of them, we measure a maximum strain amplitude larger than 1 nstrain in a visual inspection of strainmeter data, which is considered large enough to provide a sufficient signal-to-noise ratio in the DAS data. For 19 of these events, DAS data is available and allows an analysis. The largest amplitudes are found in the surface wave train. For body waves, the horizontal strain amplitude decreases as the ray incidence becomes steeper.
The table details the backazimuth (BAZ) and epicentral distance () that are given with respect to BFO (N, E), as well as the maximum strain amplitude associated with surface wave signals, . The latter is visually read from the strainmeter data and should be understood as a proxy of the actual maximum signal amplitude in the analysis.
S2.2 Strain waveforms
In Figs. S3 to S25 we display waveforms of the analyzed data. All signals are consistently filtered with Butterworth high-pass (0.05 Hz, 4th order) and low-pass (0.1 Hz, 4th order) filters. Where STS-2 data is shown, the seismometer response is equalized to the response of the strain signals. Traces are shifted vertically for better visibility.
The DAS signals from the tight-buffered cable (green) and the loose-tube cable (blue) are taken with a gauge length of 50 m at a location in the center of the ‘Anton Gang’ and the ‘Vorstollen’, respectively (see Table S1 and Figs. S1 and S2). Linear strain in azimuth NE of ‘Anton Gang’ and azimuth NE of ‘Vorstollen’ is obtained from the BFO strainmeter array like explained in the main paper.
S2.2.1 Surface wave signals
Figures S3 to S21 detail the surface wave signals of the 19 events used in the manuscript to compare the DAS recorded ‘fiber strain’ with the strainmeter recorded ‘rock strain’. The figures focus on an ad-hoc waveform comparison of surface waves for linear strain in the direction of ‘Anton Gang’ (top) and ‘Vorstollen’ (bottom). Details about the analyzed frequency band, the event origin time and beam parameters are included in the caption.
S2.2.2 Body wave signals
For the large amplitude signals of the two main shocks (Mw 7.7 and Mw 7.6) of the Kahramanmaraş earthquake sequence, we also analyze the body wave signals. The body-wave signals of the Mw 7.7 Pazarcık earthquake are displayed for the ‘Anton Gang’ in Fig. S22 and in Fig. S23 for the ‘Vorstollen’. The surface wave signals are shown in Fig. S6. The corresponding diagrams for the Mw 7.6 Ekinözü earthquake are shown in Figs. S24, S25, and S5.
ID | Mw | origin time | region | latitude | longitude | depth | BAZ | ||
/ km | / nstrain | ||||||||
gfz2023dswx | 6.79 | 2023-02-23 00:37:39.01 | Tajikistan-Xinjiang Border Region | N | E | 10.0 | NE | 20.0 | |
gfz2023doqy | 6.34 | 2023-02-20 17:04:30.09 | Turkey | N | E | 10.0 | NE | 8.0 | |
gfz2023coos | 7.59 | 2023-02-06 10:24:49.96 | Turkey | N | E | 10.0 | NE | 130.0 | |
gfz2023cnwr | 7.69 | 2023-02-06 01:17:34.97 | Turkey | N | E | 10.0 | NE | 250.0 | |
gfz2023apzg | 7.58 | 2023-01-09 17:47:35.03 | Banda Sea | N | E | 108.1 | NE | 1.5 | |
gfz2022wxsg | 6.06 | 2022-11-23 01:08:16.89 | Turkey | N | E | 10.0 | NE | 3.0 | |
gfz2022wvyo | 7.01 | 2022-11-22 02:03:07.65 | Solomon Islands | N | E | 9.7 | NE | 2.5 | |
gfz2022wpnl | 6.83 | 2022-11-18 13:37:09.2 | Southwest of Sumatra, Indonesia | N | E | 10.0 | NE | 2.0 | |
gfz2022wcnj | 7.28 | 2022-11-11 10:48:48.57 | Tonga Islands Region | N | E | 44.5 | NE | 3.0 | |
gfz2022vpri | 6.08 | 2022-11-04 10:02:51.02 | Gulf of California | N | E | 10.0 | NE | 1.2 | |
gfz2022uoko | 6.78 | 2022-10-20 11:57:11.66 | South of Panama | N | E | 10.0 | NE | 1.5 | |
gfz2022sovf | 6.79 | 2022-09-22 06:16:08.89 | Michoacan, Mexico | N | E | 15.0 | NE | 5.0 | |
gfz2022skgc | 7.59 | 2022-09-19 18:05:08.76 | Near Coast of Michoacan, Mexico | N | E | 20.0 | NE | 19.0 | |
gfz2022shod | 6.92 | 2022-09-18 06:44:16.55 | Taiwan | N | E | 10.0 | NE | 10.0 | |
gfz2022sggk | 6.46 | 2022-09-17 13:41:20.35 | Taiwan | N | E | 10.0 | NE | 2.0 | |
gfz2022rufw | 7.51 | 2022-09-10 23:46:59.95 | Papua New Guinea Region | N | E | 100.0 | NE | 2.5 | |
gfz2022rjqx | 6.60 | 2022-09-05 04:52:19.87 | Sichuan, China | N | E | 10.0 | NE | 2.0 | |
gfz2022riez | 6.88 | 2022-09-04 09:42:20.22 | Central Mid-Atlantic Ridge | N | E | 10.0 | NE | 2.5 | |
gfz2022oogo | 6.99 | 2022-07-27 00:43:25.36 | Luzon, Philippines | N | E | 21.0 | NE | 2.5 |
Appendix S3 Additional diagrams to complement the main results
Here we present some additional diagrams to support the main findings of the study. In Fig. S26 ‘strain transfer rate’ (STR) is plotted against the backazimuth (BAZ) of the respective event. We observe no significant correlation between both parameters. Fig. S27 shows STRs against NCCs. We find no correlation either. These results highlight that the STRs predominantly depends on cable type and installation conditions.
For the main shocks (Mw 7.7 and Mw 7.6) of the Kahramanmaraş earthquake sequence on February 6th 2023, we analyze the body waves in addition to the surface waves. Fig. S28 shows STRs and Fig. S29 displays NCCs. For these large amplitude signals, we observe large signal-to-noise ratio, which implies a high waveform similarity. Hence, the NCCs generally are larger than 0.92 and are largest () for the surface wave signals recorded in the ‘Anton Gang’ with the tight-buffered cable.
The ‘strain transfer rate’ is not correlated to the normalized correlation coefficient, like shown in Fig. S27, because smaller values of NCC are due to noise in the DAS data, not in the strainmeter data. In this study, we fit strainmeter against DAS data in order to determine the STR. This approach provides reasonable STR, even for heavily disturbed DAS data. This is particularly the case in the ‘Vorstollen’, with the loose-tube cable, for which NCC are globally small. Among the 19 events, illustrative examples of the situation are shown in Figs. S12, S8, S9, and S13 (bottom panels),
In the linear regression the noise in the regressor causes a bias for the regression coefficients to smaller values. If we use the DAS data as regressors in fitting the DAS waveform to the strainmeter data, we find a correlation of the regression coefficients with the NCC, which is demonstrated in Fig. S30. The signal-to-noise ratio for the strainmeter data is much better in all cases, such that no bias is found in Fig. S27, where the strainmeter data is used as regressor.
Fig. S31 displays a plot of NCC against the maximum strain amplitude. We observe a clear trend where smaller strain amplitudes are associated with lower NCC values. For smaller strain amplitudes the noise in the DAS data takes a larger fraction of the signal energy. The effect is stronger for the loose-tube cable than for the tight-buffered cable and is strongest in the ‘Vorstollen’, where cables are not embedded in sand.
Appendix S4 Plane wave strain from seismometer recordings
In this study, we fit waveforms of ‘rock strain’ recorded by the array of Invar-wire strainmeters at BFO (Zürn et al., 2015) to ‘fiber strain’ recorded by DAS. The regression coefficient is the ‘strain transfer rate’ (STR), a measure of how much of the strain amplitude is picked up by the DAS fibers. Strainmeters are rare installations and in the absence of these types of sensors studies of DAS recorded strain occasionally compare the signals with strain simulated from particle velocity recorded by seismometers, or convert DAS signals to equivalent particle velocity recordings. This conversion is possible in cases where the waveform represents a plane wave of known phase velocity and azimuth of propagation.
In section S4.1 ‘Scaling rule for plane waves’ we summarize the theory behind this conversion. The final conversion rule is given by Eq. (S6) and is valid for a single, non-dispersive plane wave.
When using surface wave-data, which provide superior signal-to-noise ratio because of their large amplitudes (compared to body-wave signals) the limitations of the plane-wave approximation become significant. The surface waves are dispersive and do not propagate with a single phase-velocity and Love- and Rayleigh-waves are inseparably superimposed. Due to their propagation along the surface of the heterogeneous Earth’s crust, they have non-plane character and contain significant components, which do not propagate along the backazimuth (BAZ) great circle. These limitation become most obvious for BAZ being perpendicular to the azimuth of measured linear strain. The plane wave does not strain the material perpendicular to the direction of propagation and value of the cosine in Eq. (S6) necessarily vanishes. The recorded signals then entirely are composed by the non-plane components of the wave, not propagating along the BAZ great circle. The problem becomes already apparent for BAZ being close to perpendicular to strain azimuth. The seismometer recorded particle velocity contains as well these non-plane components, which are at variance with the expected polarization of the plane wave. Thus the scaling in Eq. (S6) is not valid for them. We discuss the consequences of this on the regression coefficients in section S4.2 ‘Regression with respect to seismometer data’. In section S4.3 ‘Comparison of seismometer and strainmeter data’ we discuss a direct comparison of strainmeter and seismometer data.
S4.1 Scaling rule for plane waves
For signals of single non-dispersive plane waves, strain can be simulated from particle velocity recorded by a seismometer. Consider that ground deformation is due to a non-dispersive plane wave, then particle displacement at location and time is
where is the slowness-vector of the plane wave, defines the polarization and is the shape of the wave (d’Alembert’s solution to the wave equation), which propagates in direction of the slowness vector.
The linear strain in -direction for this wave is
where is the derivative of with respect to and , , and are the -components of the slowness vector , of , and of , respectively. Likewise the -component of particle velocity is
Hence
If the -direction is the horizontal direction along the ‘Anton Gang’ or the ‘Vorstollen’, given by azimuth and is the horizontal component of the plane wave slowness for propagation in a 1D structure, as can be derived by ray-tracing, then
where is the backazimuth of the source. In this way we estimate the linear strain in -direction
from the particle velocity as recorded by the broad-band seismometer.
S4.2 Regression with respect to seismometer data
Fig. S32 compares the regression coefficients measured by fitting strainmeter signals to DAS signals (filled symbols), with those obtained by fitting scaled seismometer data to DAS signals (open symbols). The seismometer data is scaled according to Eq. (S6) to represent linear strain in the direction of the DAS fiber in cases where the waves propagate along the BAZ great circle with phase slowness . This value corresponds to phase velocity of , which is the value for Rayleigh waves at 0.05 Hz (the lower end of the investigated frequency band) as shown for Southern Germany by Friederich and Huang (1996). Love-waves propagate with a higher phase velocity. For both wave types phase velocity decreases with increasing frequency, such that the chosen value lies within the range to be expected in the observed wave trains.
The regression coefficients obtained by fitting the strainmeter waveform against the DAS waveform are the STRs, as displayed in Fig. S26. They are reproduced for reference. The regression coefficients for fitting the scaled seismometer waveform against the DAS waveform show a considerably stronger scatter, not only for cases where BAZ is close to perpendicular to the strain azimuth.
Cases where BAZ is close to NE or NE have a great-circle propagation direction almost perpendicular to the azimuth of the ‘Anton Gang’ (NE). For the Vorstollen (NE), the same applies when BAZ is close to NE or NE. Most of these cases are illustrated in Fig. S32. The coefficients are significantly outside reasonable limits, with some even being negative. The scatter is more substantial for the ‘Anton Gang’, where we find several cases with BAZ almost perpendicular to the strain azimuth. For the ‘Vorstollen’ the study includes a case with BAZ near NE, but none with BAZ near NE. In consequence the overall scatter is less strong.
We also observe a systematic shift between coefficients obtained for regression with respect to strainmeter data compared to regression with respect to seismometer data. This might indicate that dominating signal propagates with a higher phase velocity than considered and that the value for the slowness in Eq. (S6) might be too large. It is, however, disputable, whether the entire difference of coefficients (a factor of 2 in some cases) in the ‘Vorstollen’ can be attributed to the wrong value of phase slowness, though there is no other obvious reason. The P-wave signals in Figs. S22, S23, S24, and S25 show an amplitude mismatch of a similar size. The mismatch varies along the P-wave-train and cannot be expressed by a single factor.
Hence, the issue is not limited to the regression coefficients. The DAS waveforms also show greater similarity to the strainmeter waveforms than to the seismometer waveforms. This is shown in Fig. S33, where the values of NCCs are plotted against backazimuth. The NNC measured by comparing DAS data with respect to seismometer data are generally smaller than for DAS with respect to the strainmeter data. For BAZ nearly perpendicular to the fiber optic cable azimuth, some of the signals are even anti-correlated (Fig. S33). This results from the change in the sign of the cosine in Eq. (S6).
The dissimilarity is attributed to the non-plane components of the waves, rather than to instrumental noise or signal-to-noise ratio (SNR) issues. Both the strainmeter and seismometer data exhibit significantly better SNR than the DAS data, with the seismometer having the highest SNR among the instruments compared. This can be seen in Figs. S22, S24, S23, and S25, where the signal level prior to the P-wave onset is least for the seismometer data, when compared to the P-wave amplitude.
S4.3 Comparison of seismometer and strainmeter data
To evaluate how well strain from particle velocity recorded by the seismometer can represent ‘rock strain’, we perform a direct comparison between the strain waveforms from the strainmeter array and those derived from the particle velocity recorded by the seismometer. Fig. S34 shows the results of the comparison. If the seismometer signal scaled by Eq. (S6) would be a good representation of ‘rock strain’, all these coefficients would equal 1. This takes place in only very few cases. If the mismatch would be due to an amplitude factor only, the coefficients for fitting the strainmeter data to the seismometer data would be the reciprocals of the coefficients computed when fitting seismometer data to strainmeter data. This is not the case, due to a considerable waveform mismatch caused by non-plane components of the surface waves.